# A meshless method for computational electromagnetics with improved dispersion properties

### v1

@@ 5 Generalising the FDTD to a meshless setting
Having the tools of numerical analysis on scattered nodes behind us, we can now proceed to implementing a variant of the FDTD in a meshless setting.
Firstly, the previously mentioned shortcomings of FDTD arise due to the grid discretisation in space. Time direction has a trivial geometry and poses no issue, so we will keep the staggering of fields in time and the same time evolution method - the interleaved leapfrog.
The problematic part is the spatial discretisation, which will be given by scattered nodes. We have decided to discard the idea of staggering the $\vec{H}$ and $\vec{E}$ fields at different points, as this could become nontrivial in complex settings (even more so in higher dimensions). Instead, we simply take the FDTD update equations (5) and replace the spatial derivatives (central differences) by their suitable meshless substitutes:
$H_{x}(x_{i},y_{i},n+1)=H_{x}(x_{i},y_{i},n)-\frac{\Delta t}{\mu_{0}}D^{\partial_{y}}(E_{z}(x_{i},y_{i},n)),$ (16)
$H_{y}(x_{i},y_{i},n+1)=H_{y}(x_{i},y_{i},n)+\frac{\Delta t}{\mu_{0}}D^{\partial_{x}}(E_{z}(x_{i},y_{i},n)),$
@@
In the first approach, we will calculate the differentiation weights using RBF-VFD, where the finite difference stencils in question are simply the central differences. This approach seems to be as close to the Yee grid as we can get to in the meshless setting. We have opted for $\delta=0.5h$ as our virtual stencil spacing of choice, which is within the previously mentioned guidelines and mirrors the spacing of the original Yee grid, where the $\vec{E}$ and $\vec{H}$ are also offset by the same amount.
In the second approach we will apply RBF-FD to obtain the differentiation weights, which directly generalises the usual FDTD [15].
In order to match the order of FDTD, we calculate the derivative weights using $m=2$.
The above scheme remains explicit and thus efficient, as the differentiation weights are only computed once at the beginning of the simulation and can be reused. We gain additional flexibility of the method due to the meshless nature of the differential operators.
However, our attempt at a more geometrically flexible version of the FDTD comes at a cost in terms of the method stability.
In FDTD it can be shown that the method is stable if the Courant number satisfies $S_{c}:=\frac{c_{0}\Delta t}{\Delta s}\leq S_{c}^{\mathrm{max}}$, where $S_{c}^{\mathrm{max}}$ equals $\frac{1}{\sqrt{2}}$ for the two-dimensional case with similar stability conditions on the Courant number available also for other regular node layouts [12]. However, in a meshless setting, the nodes are scattered and such an estimate is not easily reachable. We have attempted to stabilise the method by computing at an increasingly lower $S_{c}$ with no success [^6] For the forthcoming analyses, we have set $S_{c}=(10\sqrt{2})^{-1}$., hinting at an inherent instability of the method. Refer to the Appendix for an illustration of why instability may be expected. To remedy these problems and make our proposed methods useful, a stabilisation procedure is needed.
@@ 5.1 Stabilising the method
Instabilities in explicit numerical schemes are a well-known problem and subject to a plethora of research in the scientific community. Many different approaches to stabilisation have been proposed over the years, such as penalty schemes or upwinding, to name a few.
We have opted for an approach that has been gaining popularity in the meshless community lately: Hyperviscosity (HV) [10]. It can be effective and is in principle very easy to implement, as it amounts to simply adding a higher order derivative term to our system:
@@
where $\Delta$ is the Laplace operator. For further intuition on how HV stabilises a given method, an interested reader may refer to [23, 10].
What remains is a determination of parameters $\gamma_{i},\alpha_{i}$ that should be set such that the HV term stabilises the solution, while not affecting the physically relevant solution components too much. We first perform some simplifications that turned out to work well. We set:
@@
Additionally, we follow the hyperviscosity guidelines described in [23] that set $\gamma=c\ (-1)^{\alpha-1}h^{2\alpha}$, for some constant $c$. We will discretise the newly introduced hyperviscosity operator $\gamma_{i}\Delta^{\alpha}$ using RBF-FD, resulting in modified update equations of the stabilised method:
$H_{x}(x_{i},y_{i},n+1)=H_{x}(x_{i},y_{i},n)-\frac{\Delta t}{\mu_{0}}\left(D^{\partial_{y}}(E_{z}(x_{i},y_{i},n))-\gamma D^{\Delta^{\alpha}}H_{x}(x_{i},y_{i},n)\right),$
$H_{y}(x_{i},y_{i},n+1)=H_{y}(x_{i},y_{i},n)+\frac{\Delta t}{\mu_{0}}\left(D^{\partial_{x}}(E_{z}(x_{i},y_{i},n))+\gamma D^{\Delta^{\alpha}}H_{y}(x_{i},y_{i},n)\right),$ (19)
$E_{z}(x_{i},y_{i},n+1)=E_{z}(x_{i},y_{i},n)+\frac{\Delta t}{\epsilon_{0}}\left(D^{\partial_{x}}H_{y}(x_{i},y_{i},n+1)-D^{\partial_{y}}H_{x}(x_{i},y_{i},n+1)+\gamma D^{\Delta^{\alpha}}E_{z}(x_{i},y_{i},n)\right).$
Note that, for simplicity, we will use RBF-FD for the HV part even if the $\partial_{x},\partial_{y}$ operators are discretised using RBF-VFD. The rationale is that the HV term is merely a stabilisation procedure and should not be responsible for other properties of the method. Following also the guidelines presented in [32], we choose $\varphi(r)=r^{2\alpha+1}$, $m=2\alpha$ and $n=2M_{m}+1$ for the RBF-FD discretisation of the HV operator.
The values of the parameters $\alpha$ and $c$ are determined by a parameter sweep - we solve a given problem with some chosen parameters and observe its stability properties.
@@
$U(t)$ $\propto\int_{\Omega}\|\vec{E}(t)\|^{2}+\eta_{0}^{2}\|\vec{H}(t)\|^{2}\textup{d}\Omega$
@@
where the integral was approximated with a sum similarly as in [31]. Due to staggering in time, we do not have access to the $\vec{E}$ and $\vec{H}$ fields at the same time points. In order to calculate the energy at an integer multiple of $\Delta t$, we must average the $\vec{H}$ fields (which are otherwise only defined at odd multiples of $\Delta t/2$). We should emphasize that an analytical solution of the problem is not needed to compute $U(t)$. In practice, we can observe the relative energy $u(t)=U(t)/U(0)$, which should remain approximately $1$ for a conservative problem setup.
@@
Figure 5: The behaviour of relative energy $u(t)$ over time for the considered test case for different hyperviscosity parameters $\alpha$ and $c$.
The time dependence of $u(t)$ for different stabilisation parameters is displayed in Figure 5. We can observe that $c$ has a major effect on the stability of the method and that a well-chosen value of $c$ and $\alpha$ leads to an approximate conservation of energy throughout the simulation. A more complete picture of the effect of $c$ on stability can be seen in Figure 6, where the energy after $10000$ timesteps with respect to $c$ is shown. Here, a common pattern in HV can be observed: There exists an interval of suitable values of $c$, for which the solution is stable. For $c$ outside this interval, the solution diverges. Note that stable schemes may also be dissipative, which in our case happens at $\alpha=1$ and is undesirable. In more complex settings the interval of $c$, for which the method is stable, can be very narrow and dissipation commonly arises, even in $\alpha>1$ case. For that reason, complex settings may require a very fine parameter sweep and for practical reasons, we may to resort to the bisection approach to efficiently calculate a suitable value of $c$.
For the remainder of this section, we have settled on $\alpha=2$ and $c=10^{-4}$, as it is a stable and non-dissipative combination. Generally, we prefer a lower $\alpha$ if possible, as the stencil size and therefore the computational complexity of the method grows with $\alpha$.
@@ 5.2 Convergence of the solution
where we have added the superscript $a$ to denote the analytical solution and differentiate from the numerically obtained solution. This allows us to verify that our stabilised method is indeed convergent. We check this by reducing the spatial and time discretisation, while keeping the Courant number $S_{c}=\frac{1}{10\sqrt{2}}$ constant. To study the convergence we will define the following two $\ell^{\infty}$ errors:
@@
The results are displayed in Figure 7, where we can see that both methods are convergent with similar errors behaviour for both methods as $h\to 0$. This dependence maintains its shape also if we consider slightly different discretisations by modifying the random seed in our node generation algorithm [25]. Ideally, a convergence rate of $\propto h$ is expected to occur, however, a slightly slower convergence is not surprising due to hyperviscosity.
Figure 7: $\|e\|^{H}_{\infty}$ and $\|e\|^{E}_{\infty}$ errors with respect to the discretisation distance $h$.
To summarise, we now have a stable method that can operate on scattered nodes. With this we have addressed the first shortcoming of the FDTD and we now turn to the second - anisotropical behaviour of the dispersion relation.
@@ 6 Dispersion
Having developed a working meshless algorithm, we now turn to its dispersion properties.
@@
To evaluate the magnitude of dispersion of our methods, we consider a similar test case as in Section 5.1, except that we use $\sigma=2$ - a very narrow Gaussian that is difficult to resolve at $h=1$ and will push the method to its limits inducing visible effects of dispersion. As the setup is now different, the previously calculated HV parameters are not suitable and we recompute them in the same manner as before. As the newly considered setup is much more unstable, a naive parameter sweep is slow and we have computed the parameters using the bisection approach. We have settled for $\alpha=2$ and $c\approx 10^{-4.76}$ for the RBF-FD and $c\approx 10^{-5.02}$ for RBF-VFD, which maintained $u(t)\in[0.99,1.01]$ even after $10000$ timesteps.
@@
To analyse dispersion we will turn to Fourier analysis, which works well on uniformly spaced nodes. For that reason, we perform an extra step when discretising $0,200]^{2} -$ first, we discretise the line $y=100$ with equispaced nodes $\{\mathbf{x}^{e}_{i}\}_{i}$, at a distance of $h=1$ apart. These are then used as seed nodes in the same discretisation algorithm to obtain the complete discretisation of $0,200]^{2}$, effectively resulting in a scattered node layout, where some points are guaranteed to be equispaced on a line.
@@
Figure 8 shows the snapshot of the simulation after $2820$ timesteps (approximately equal to one passage of the wavefront across the whole domain) for different method parameters, where several observations can be made. First, we can see that increasing the stencil size [^7] Note that this refers to the stencil size of the original, unstabilised parts of the update equations. For stabilisation, we continue to adhere to the previous guidelines. improves the dispersion properties of the solution. In fact, for this particular case, the smallest recommended stencil size $n=13$ is very sensitive to the properties of the discretisation, resulting in numerical artifacts in the solution. We do not yet fully understand the reasoning behind this behaviour, although some ideas on how to analyse it will be mentioned in the conclusion. A second thing to observe is that, perhaps surprisingly, dispersion seems much more prominent in the RBF-VFD approach compared to the RBF-FD.
Figure 8: Snapshot of the solution for different stencil sizes and methods.
@@
To put our observations in a formal setting, let us analyse the spectral picture of the solution. Running the simulation as before, we sample the solution values $E_{z}(\mathbf{x}^{e}_{i},n\Delta t)$ and perform a two-dimensional Fourier transform to obtain the corresponding values in frequency space $E_{z}(k,\omega)$, where the allowed discrete values for $k$ are $k_{i}=\frac{2\pi i}{N_{x}}$, for $0\leq i<N_{x}$, with $N_{x}=200$ (number of points on the discretised line) and similarly $\omega_{j}=\frac{2\pi j}{15N_{t}}$ with $0\leq j<N_{t}$ with $N_{t}=188$ (running the simulation for $2820$ timesteps and saving every 15th snapshot). $k$ are given in units of $h^{-1}]$, while the units of $\omega$ are $(\Delta t)^{-1}]$.
@@
Before transforming, Hann window was applied in the time direction to prevent spectral leakage due to non-periodicity (in time) of the solutions. The results can be seen in Figure 9, where the absolute value of the individual spectral components $E_{z}(k_{i},\omega_{j})$ is displayed. Physically, we would expect all the modes to propagate at the same speed, that is, that the magnitude of $E_{z}(k,\omega)$ is large only when $\omega=c_{0}k$, where $c_{0}=\frac{1}{10\sqrt{2}}$ in our units. A black, dashed line representing this physical dispersion relation is also displayed in the same Figure. We can clearly see the different behaviour of the cases with the smallest stencil size ($n=13$), where the linear dispersion relation breaks down already at a relatively low frequency. This is greatly improved as stencil size is increased (note the logarithmic scale - the off-diagonal components are multiple orders of magnitude smaller). As already hinted by our previous simulation, we can observe that the RBF-VFD approach further deviates from the expected dispersion relation compared to the RBF-FD.
@@
For the purposes of visualisation, we would like to summarise the amount of error in the numerical dispersion with a single quantity. We first define a numerical propagation velocity $c(\omega_{j})$ by selecting for each discrete frequency $\omega_{j}$ the corresponding wave vector $k_{i}$ that is most represented:
$i(j)$ $\mathrel{\mathop{:}}=\mathrm{argmax}_{0\leq i<N_{x}}|E(k_{i},\omega_{j})|$ (23)
$c(\omega_{j})$ $\mathrel{\mathop{:}}=\frac{\omega_{j}}{k_{i(j)}}$
@@
$\mathbf{err}(\omega_{j})\mathrel{\mathop{:}}=\left|c(\omega_{j})-\frac{1}{10\sqrt{2}}\right|$ (24)
As our quantity, summarising the amount of deviation from the numerical dispersion relation, we simply take the sligthly modified $\ell^{1}$ norm of $\mathbf{err}$:
@@
where we have only summed up to $j=45$, as higher values are not visibly present in the spectrum and would distort the results. This quantity is displayed in Figure 10 with respect to the stencil size and matches our previous observations - the error is large at small stencil sizes and is reduced as the stencil size is increased. From this point onward, we will work with $n=60$ as our stencil size of choice for both methods.
Figure 10: Stencil size dependence of the chosen dispersion error.
Having delved into the dispersive properties of our presented methods it is time to introduce angular dependence and compare it to the FDTD. For this, we modify our setup yet again - we again start with $0,200]^{2}$ and a line $y=100$, however, we now rotate it by a chosen angle $\phi$, and afterwards complete the discretisation, again using the (now rotated) line as seed nodes. The initial condition is the same pulse wave as before, moving along the square (in the $\phi$ direction):
@@
Such a setup allows us to maintain the periodic boundary condition and perform the same Fourier analysis as before, the difference being that the pulse now propagates at an angle $\phi$ with respect to the $xy$ coordinate system.
The results of the simulation for two extreme angles $\phi=0$ and $\phi=\frac{\pi}{4}$ are seen in Figure 11. Note that, while for the $\phi=0$ case, we can implement the FDTD on $0,200]^{2}$ as before, the rotated case was implemented on $0,144\sqrt{2}]^{2}$, rotated by $\frac{\pi}{4}$, as we cannot make the grid periodic otherwise. Qualitative results should be independent of the slight change in the domain size, as the width of the pulse remains the same. We can observe that the RBF-FD case is not affected by the orientation, which is what we initially hoped for, as the stencil has no directional bias. This directional independence was not a-priori obvious, however: In the RBF-FD approach we still compute the derivatives in the $x$ and $y$ direction separately and do not manually account for the fact that the curl operator is rotationally invariant.
On the contrary, the dispersion properties of the FDTD are expectedly highly angle-dependent. Interestingly, the same holds for RBF-VFD, implying that having a stencil with no directional bias is not sufficient for isotropy and the anisotropy can arise from the manner in which differentiation weights are computed.
Figure 11: Snapshots of the simulation after one passage through the domain for the two extremal angles.
We conclude the section by plotting $\|\mathrm{err}\|_{1}(\phi)$ in Figure 12, where the RBF-FD and RBF-VFD curves were computed in the same manner as before, while for the FDTD we calculated the error from Equation (6). The latter plot again shows that the RBF-FD approach exhibits a much lower degree of dispersion anisotropy, compared to the RBF-VFD and the FDTD approaches, which behave similarly.
Figure 12: Dispersion error dependence on the propagation angle $\phi$.

### v2

@@ 5 Generalising the FDTD to a meshless setting
Having the tools of numerical analysis on scattered nodes behind us, we can now proceed and implement a variant of the FDTD in a meshless setting.
Firstly, the previously mentioned shortcomings of FDTD arise due to a regular discretisation in space. Time direction has trivial geometry and poses no issues, so we will keep the staggering of fields in time and the same time evolution method as in the FDTD – the interleaved leapfrog.
The problematic part is the spatial discretisation, which will be given by scattered nodes. As mentioned in the introduction, we have decided to discard the idea of staggering the $\vec{H}$ and $\vec{E}$ fields at different points. Instead, we simply take the FDTD update equations (5) and replace the spatial derivatives (central differences) by their suitable meshless substitutes:
$H_{x}(x_{i},y_{i},n+1)=H_{x}(x_{i},y_{i},n)-\frac{\Delta t}{\mu_{0}}D^{\partial_{y}}(E_{z}(x_{i},y_{i},n)),$
$H_{y}(x_{i},y_{i},n+1)=H_{y}(x_{i},y_{i},n)+\frac{\Delta t}{\mu_{0}}D^{\partial_{x}}(E_{z}(x_{i},y_{i},n)),$ (16)
@@
In the first approach, we will calculate the differentiation weights using RBF-VFD, where the finite difference stencils in question are simply the central differences. This approach seems to be as close to the Yee grid as we can get to in a meshless setting. We have opted for $\delta=0.5h$ as our virtual stencil spacing of choice, which is within the previously mentioned guidelines and mirrors the spacing of the original Yee grid, where the $\vec{E}$ and $\vec{H}$ are also offset by the same amount. In the second approach we will apply RBF-FD to obtain the differentiation weights, which directly generalises the usual FDTD [15]. In order to match the order of FDTD, we calculate the derivative weights using $m=2$ in both approaches.
The above scheme remains explicit and thus efficient, as the differentiation weights are only computed once at the beginning of the simulation and can be reused. We gain additional flexibility of the method due to the meshless nature of the differential operators. However, our attempt at a more geometrically flexible version of the FDTD comes at a cost in terms of the method stability. As mentioned, in FDTD it can be shown that the method is stable if the Courant number satisfies $S_{c}:=\frac{c_{0}\Delta t}{\Delta s}\leq S_{c}^{\mathrm{max}}$, where $S_{c}^{\mathrm{max}}$ equals $\frac{1}{\sqrt{2}}$ for the two-dimensional case with similar stability conditions available also for other regular node layouts [12]. However, in a meshless setting, the nodes are scattered and such an estimate is not easily reachable. We have attempted to stabilise the method by computing at an increasingly lower $S_{c}$ with no success [^7] For the forthcoming analyses, we have set $S_{c}=(10\sqrt{2})^{-1}$., hinting at an inherent instability of the method (refer to the Appendix for an illustration of why instability may be expected). To remedy these problems and make our proposed methods useful, a stabilisation procedure is needed.
@@ 5.1 Stabilising the method
Instabilities in explicit numerical schemes are a well-known problem and subject to a plethora of research in the scientific community. Several different approaches to stabilisation have been proposed over the years, such as penalty schemes [33] or upwinding [10], to name a few.
We have opted for an approach that has been gaining popularity in the meshless community lately: yperviscosity (HV) [9]. It can be effective and is in principle very easy to implement, as it amounts to simply adding a higher order derivative term to our system:
@@
where $\Delta^{\alpha}$ is an iterated Laplace operator. For further intuition on how HV stabilises a given method, we refer an interested reader to [24, 9].
What remains is to determine parameters $\gamma_{i},\alpha_{i}$, which should be such that the HV term stabilises the solution, while minimally affecting the physically relevant solution components. We first perform some simplifications that turned out to work well. We set:
@@
Additionally, we follow the hyperviscosity guidelines described in [24] and set $\gamma=c\ (-1)^{\alpha-1}h^{2\alpha}$, for some $c>0$. We will discretise the newly introduced hyperviscosity operator $\Delta^{\alpha}$ using RBF-FD, resulting in modified update equations of the stabilised method:
$H_{x}(x_{i},y_{i},n+1)=H_{x}(x_{i},y_{i},n)-\frac{\Delta t}{\mu_{0}}\big($ $D^{\partial_{y}}(E_{z}(x_{i},y_{i},n))-\gamma D^{\Delta^{\alpha}}H_{x}(x_{i},y_{i},n)\big),$
$H_{y}(x_{i},y_{i},n+1)=H_{y}(x_{i},y_{i},n)+\frac{\Delta t}{\mu_{0}}\big($ $D^{\partial_{x}}(E_{z}(x_{i},y_{i},n))+\gamma D^{\Delta^{\alpha}}H_{y}(x_{i},y_{i},n)\big),$
$E_{z}(x_{i},y_{i},n+1)=E_{z}(x_{i},y_{i},n)+\frac{\Delta t}{\epsilon_{0}}\big($ $D^{\partial_{x}}H_{y}(x_{i},y_{i},n+1)-D^{\partial_{y}}H_{x}(x_{i},y_{i},n+1)\ +$
$+\ \gamma$ $D^{\Delta^{\alpha}}E_{z}(x_{i},y_{i},n)\big).$ (19)
Note that, for simplicity, RBF-FD is used for HV terms even if the $\partial_{x},\partial_{y}$ operators are discretised using RBF-VFD. The rationale behind this decision is that HV terms are merely a stabilisation procedure and should not be responsible for other properties of the method. Following also the guidelines presented in [34], we choose $\varphi_{\mathrm{HV}}(r)=r^{2\alpha+1}$, $m_{\mathrm{HV}}=2\alpha$ and $n_{\mathrm{HV}}=2M_{m_{\mathrm{HV}}}+1$ for the RBF-FD discretisation of the HV operator.
The values of parameters $\alpha$ and $c$ are determined by a parameter sweep – we solve a given problem with some chosen parameters and observe its stability properties.
@@
$U(t)$ $\propto\int_{\Omega}\|\vec{E}(t)\|^{2}+\eta_{0}^{2}\|\vec{H}(t)\|^{2}\textup{d}\Omega\approx$ (20)
@@
where the integral was approximated with a sum similarly as in [32]. Due to staggering in time, we do not have access to $\vec{E}$ and $\vec{H}$ fields at the same time points. In order to calculate the energy at an integer multiple of $\Delta t$, we must average the $\vec{H}$ fields (which are otherwise only defined at odd multiples of $\Delta t/2$). In practice, we track the relative energy $u(t)=U(t)/U(0)$, which should remain approximately $1$ for a conservative problem setup. We should emphasize that no information on analytical solution of the problem is needed to compute $u(t)$.
@@
Figure 5: Behaviour of relative energy $u$ over time for the considered test case for different hyperviscosity parameters $\alpha$ and $c$.
The time dependence of $u(t)$ for different stabilisation parameters is displayed in Figure 5, where the problem was solved with RBF-VFD. We can observe that $c$ has a major effect on the stability of the method and that a well-chosen value of $c$ and $\alpha$ leads to an approximate conservation of energy throughout the simulation. A more complete picture of the effect of $c$ on stability can be seen in Figure 6, where relative energy after $10000$ timesteps with respect to $c$ is shown. Here, a common pattern in HV can be observed: There exists an interval of suitable values of $c$, for which the solution is stable. For $c$ outside this interval, the solution diverges. Note that stable schemes may also be dissipative, which in our case happens for $\alpha=1$ and is undesirable. In more complex settings dissipation commonly arises, even in $\alpha>1$ case. In that case a very fine parameter sweep would be required to find a stable value of $c$ with little dissipation and for practical reasons, we resort to the bisection approach as explained in [20].
For the remainder of this section, we have settled on $\alpha=2$ and $c=10^{-4}$ as it is a stable combination with little dissipation. Generally, we prefer a lower $\alpha$ if possible, as the stencil size and therefore the computational complexity of the method grows with $\alpha$.
@@ 5.2 Convergence of the solution
where we have added the superscript $a$ to denote the analytical solution and differentiate it from the numerically obtained solution. This allows us to verify that our stabilised method is indeed convergent. We check this by reducing discretisation distances in space and time, while keeping the Courant number $S_{c}=\frac{1}{10\sqrt{2}}$ constant. To study convergence we will define the following two $\ell^{\infty}$ errors:
@@
Results are displayed in Figure 7, where we can see that both methods are convergent with similar error behaviour as $h\to 0$. Ideally, a convergence rate of $\propto h$ is expected to occur, however, a slightly slower convergence is not surprising due to the presence of hyperviscosity.
Figure 7: $\|e\|^{H}_{\infty}$ and $\|e\|^{E}_{\infty}$ errors with respect to the discretisation distance $h$. Errors were computed at time $c_{0}t=5\sqrt{2}$.
To summarise, we now have a stable method that can operate on scattered nodes. With this we have addressed the first shortcoming of the FDTD and we now turn to the second - anisotropic dispersion relation.
@@ 6 Dispersion
Having developed a working meshless algorithm, we now turn to its dispersion properties. To evaluate the magnitude of dispersion of our methods, we consider a similar test case as in Section 5.1, except that we use $\sigma=2$ - a very narrow Gaussian that is difficult to resolve at $h=1$ and will push the method to its limits inducing visible effects of dispersion. As the setup is now different, the previously calculated HV parameters are not suitable and we recompute them. Since the newly considered setup is much more unstable, a naive parameter sweep is slow and we have computed the parameters using the previously mentioned bisection approach. We have settled for $\alpha=2$ and $c\approx 10^{-4.76}$ for the RBF-FD and $c\approx 10^{-5.02}$ for RBF-VFD, which maintained $u(t)\in[0.99,1.01]$ even after $10000$ timesteps.
@@
To analyse dispersion we will turn to Fourier analysis, which works well on uniformly spaced nodes. For that reason, we perform an extra step when discretising $\Omega=[0,200]^{2}$ - first, we discretise the line $y=100$ with equispaced nodes $\{\mathbf{x}^{e}_{i}\}_{i}$, at a distance of $h=1$ apart. These are then used as seed nodes in the same discretisation algorithm to obtain a complete discretisation of $\Omega$, effectively resulting in a scattered node layout, where some points are guaranteed to be equispaced on a line.
@@
Figure 8 shows the snapshot of the simulation after $2820$ timesteps (approximately equal to one passage of the wavefront across the whole domain) for different method parameters, where several important observations can be made. First, we can see that increasing the stencil size [^8] Note that this refers to the stencil size of the original parts of the update equations. For HV terms, we continue to adhere to the previous guidelines. improves the dispersion properties of the solution. In fact, for this particular case, the smallest recommended stencil size $n=13$ contains numerical artefacts in the solution. We do not yet fully understand the reasoning behind this behaviour, although some ideas on how to analyse it will be mentioned in the conclusion. A second thing to observe is that, perhaps surprisingly, dispersion seems much more prominent in the RBF-VFD approach compared to the RBF-FD.
Figure 8: Snapshot of the solution at $c_{0}t=282$ for different stencil sizes and methods.
@@
To put our observations in a quantitative setting, let us analyse the spectral picture of the solution. Running the simulation as before, we sample the solution values $E_{z}(\mathbf{x}^{e}_{i},n\Delta t)$ and perform a two-dimensional Fourier transform to obtain the corresponding values in frequency space $E_{z}(k,\omega)$, where the allowed discrete values for $k$ are $k_{i}=\frac{2\pi i}{N_{x}}$, for $0\leq i<N_{x}$, with $N_{x}=200$ (number of points on the discretised line) and similarly $\omega_{j}=\frac{2\pi j}{15N_{t}}$ with $0\leq j<N_{t}$ with $N_{t}=188$ (running the simulation for $2820$ timesteps and saving every 15th snapshot).
@@
Before transforming, Hann window was applied in the time direction to prevent spectral leakage due to non-periodicity (in time) of the solutions. The results can be seen in Figure 9, where the absolute value of the individual spectral components $E_{z}(k_{i},\omega_{j})$ is displayed. Physically, we would expect all the modes to propagate at the same speed, that is, that the magnitude of $E_{z}(k,\omega)$ is large only when $\omega=c_{0}k$, where $c_{0}=\frac{1}{10\sqrt{2}}$ in our units. A black, dashed line representing this physical dispersion relation is also displayed in the same Figure. We can clearly see the different behaviour of the cases with the smallest stencil size ($n=13$), where the linear dispersion relation breaks down already at a relatively low frequency. This is greatly improved as stencil size is increased (note the logarithmic scale – the off-diagonal components are several orders of magnitude smaller). As already hinted by our previous simulation, we can observe that the RBF-VFD approach further deviates from the expected dispersion relation compared to the RBF-FD.
@@
For the purposes of visualisation, we would like to summarise the amount of numerical dispersion with a single quantity. We first define a numerical propagation velocity $c(\omega_{j})$ by selecting for each discrete frequency $\omega_{j}$ the corresponding wave vector $k_{i}$:
$i(j)$ $\mathrel{\mathop{:}}=\mathrm{argmax}_{0\leq i<N_{x}}|E(k_{i},\omega_{j})|,$ (24)
$c(\omega_{j})$ $\mathrel{\mathop{:}}=\frac{\omega_{j}}{k_{i(j)}}.$
@@
$\mathbf{err}(\omega_{j})\mathrel{\mathop{:}}=\left|c(\omega_{j})-\frac{1}{10\sqrt{2}}\right|.$ (25)
As a quantity that summarises the amount of deviation from the physical numerical dispersion relation, we simply take the sligthly modified $\ell^{1}$ norm of $\mathbf{err}$:
@@
where we have only summed up to $j=45$, as higher values are not visibly present in the spectrum and would distort the results. This quantity is displayed in Figure 10 with respect to the stencil size and matches our previous observations – numerical dispersion is large at small stencil sizes and is reduced as the stencil size is increased. From this point onward, we will work with $n=60$ as our stencil size of choice for both methods.
Figure 10: Stencil size dependence of numerical dispersion.
Having delved into the dispersive properties of our presented methods it is time to introduce angular dependence and compare it to the FDTD. For this, we modify our setup yet again – we again start with $\Omega=[0,200]^{2}$ and a discretised $y=100$ line, rotate it by a chosen angle $\phi$, and afterwards complete the discretisation. The initial condition is effectively the same pulse wave as before, moving along the square (in the $\phi$ direction):
@@
where $s_{\phi}=\sin(\phi)$ and $c_{\phi}=\cos(\phi)$. Such a setup allows us to maintain periodic boundary conditions and perform the same Fourier analysis as before, the difference being that the pulse now propagates at an angle $\phi$ with respect to the $xy$ coordinate system.
Results of the simulation for two extreme angles $\phi=0$ and $\phi=\frac{\pi}{4}$ are seen in Figure 11. Note that, while for the $\phi=0$ case, we can implement the FDTD on $0,200]^{2}$ as before, the rotated case was implemented on $0,144\sqrt{2}]^{2}$, rotated by $\frac{\pi}{4}$, as we cannot make the grid periodic otherwise. Qualitative results should be independent of the slight change in the domain size, as the width of the pulse remains the same. We can observe that the RBF-FD case is not affected by the orientation, which is what we initially hoped for, as the stencil has no directional bias. This directional independence was not a-priori obvious, however: In the RBF-FD approach we still compute the derivatives in the $x$ and $y$ direction separately and do not at any point account for the fact that the curl operator is rotationally invariant.
On the contrary, the dispersion properties of the FDTD are expectedly highly angle-dependent. Interestingly, the same holds for RBF-VFD, implying that having a stencil with no directional bias is not sufficient and the anisotropy can arise from the manner in which differentiation weights are computed.
Figure 11: Snapshots of the simulation after one passage through the domain for two extremal angles.
We conclude the section by plotting $\|\mathbf{err}\|_{1}(\phi)$ in Figure 12, where the RBF-FD and RBF-VFD curves were computed in the same manner as before, while for the FDTD we calculated $\mathbf{err}$ from Equation (6). This plot again shows that the RBF-FD approach exhibits a much lower degree of dispersion anisotropy compared to the RBF-VFD and the FDTD, which behave similarly.
Figure 12: Numerical dispersion dependence on the propagation angle $\phi$.